The statement of financial position of KY Limited at 31 December 2019 shows the following. Non-current assets GHC 1,350,000; current assets GHC 140,000; ordinary share capital GHC 900,000; general reserve GHC 150,000; 10% debentures GHC 200,000; current liabilities GHC 90,000; retained earnings GHC 150,000 The profit from operations for the year was GHC 65 000 and the finance costs were GHC 20 000. What is the return on capital employed for 2019?
1
Expert's answer
2020-07-03T09:46:01-0400

Formula of Return on Capital Employed (ROCE) = Earnings before Interest and Tax (EBIT)/ Capital Employed (Total assets - current liabilities)

Earnings before interest and tax (EBIT) = GHC 65000

Total assets = 1350000+140000= GHC 1490000

Current liabilities = GHC 90000

Capital employed = 1490000 - 90000 = GHC 1400000

Return on capital employed = 65000 / 1400000 x 100 = 4.64%